• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name cisbani

    Log entry time 08:04:29 on November16,2008

    Entry number 249170

    keyword=Owl Shift Summary

    Owl Shift Summary (Nov/16/2008)

    00:00 tuning beam, problem in getting harp.

    01:30 possible solution for the Harp scan problem: reboot the iocha1 (MCC will do that)

    03:10 Start Production on empty ref. cell (-023 Torr, -019 PSIG),
    Happex run 30410, Run 3541, I=2 uA

    04:20 Some BB chamber planes tripped twice in the same run 3542

    05:00 Moved to 3He - Transv. Pol 57.8%

    05:10 Started production on 3He Tpol run 3544, I=2 uA

    06:30 Short study on origin of deadtime

    06:50 Asked for I=4 uA

    06:55 Happex run terminated; a new one started: 30411

    07:00 Production run 3550 at 4 uA and T1=100 Hz

    07:40 Asked for 8 uA